OpenStudy (jchick):

Simplify completely: the quantity 6 times x to the 4th power plus 9 times x to the 2nd power plus 12 times x all over 3 times

OpenStudy (jchick):

a. 2x3 + 9x2 + 12 b. 2x2 + 3x + 4 c. 2x5 + 3x3 + 4x2 d. 2x3 + 3x + 4

OpenStudy (jhannybean):

Can you write out your function first?

OpenStudy (jchick):

Here it is

OpenStudy (jhannybean):

Since x is common to all terms in both numerator and denominator, and 6,9, and 12 are a multiple of 3, reduce each term in the numerator by 3x

OpenStudy (jhannybean):

so \[\frac{6x^4}{3x} ~,~ \frac{9x^2}{3x} ~,~ \frac{12x}{3x}\]

OpenStudy (jchick):

I am lost

OpenStudy (jhannybean):

what is 6/3?

OpenStudy (jchick):

2

OpenStudy (jhannybean):

and \(\large x^{4-1}\)?

OpenStudy (jchick):

x^3

OpenStudy (jhannybean):

That takes care of the first part of the problem. You get \(2x^3\).

OpenStudy (jhannybean):

Now what is 9/3 and \(\large x^{2-1}\)?

OpenStudy (jchick):

3 * x^1

OpenStudy (jhannybean):

and so the second part gives you \(3x\).

OpenStudy (jhannybean):

Now for the final part: 12/3 and \(x^{1-1}\)

OpenStudy (jchick):

4x

OpenStudy (jhannybean):

\[x^{1-1} =x^0 =1\] That's just an identity you learn and get comfortable with.

OpenStudy (jhannybean):

So therefore it'd just be 4.

OpenStudy (jhannybean):

Now we just combine all our results together to get: \(2x^3+3x+4\)

OpenStudy (jchick):

That isn't one of the options

OpenStudy (jchick):

Nevermind

OpenStudy (jhannybean):

:)